OpenStudy (anonymous):

The midpoint of line UV is (5, –11). The coordinates of one endpoint are U(3, 5). Find the coordinates of endpoint V. Alright , I know how to do this on a graph , but how do I do this with a mathematical equation ? That's what it's requesting , and I don't recall being taught how to solve it off of a graph o.e Please explain steps in your answer , thank you <3

OpenStudy (anonymous):

you can use algebra or you can do it in your head

OpenStudy (anonymous):

from 3 to 5 on the \(x\) axis is two units right. go another two units, you get to 7

OpenStudy (anonymous):

Like it says in the post; I know how to do it on a graph ; which is what you explained , but it is asking for a mathematical equation to do this .

OpenStudy (anonymous):

from 5 to -11 in the \(y\) direction is 16 units down got down another 16, you are at -27

OpenStudy (anonymous):

then solve \[\frac{3+x}{2}=5\] and \[\frac{5+y}{2}=-11\]

Parth (parthkohli):

Don't use equations. Remember that the \(x\) and \(y\) coordinates in the midpoint are just the averages of both \(x\) and \(y\) coordinates.

OpenStudy (anonymous):

an equation is nice, but thinking is always a plus

OpenStudy (anonymous):

let (x,y)= (5,-11) and (x1,y1) = (3,5) we have a formula for that (2x-x1,2y-y1) substitute it and you can get the answer for the question...

Parth (parthkohli):

In other words, if two points are \((0,1)\) and \((5,6)\), then the midpoint is \(\left({0 +5 \over 2} , {1 + 6 \over 2}\right)\)

Parth (parthkohli):

And if three points are \((0,1), (5,7)\) and \((9,8)\), then your midpoint is,\[\left({0 + 5 + 9 \over 3} , {1 + 7 + 8 \over 3} \right)\]

OpenStudy (anonymous):

-I am required to use an equation ; again , like it says in the post .. -& I know how to find the midpoint , but it's not asking for the midpoint . I am provided 1 endpoint and the midpoint to solve for the other endpoint . -Answers which applied to what I am really asking didn't show all steps / explain ._.

OpenStudy (anonymous):

. . .

OpenStudy (anonymous):

Nobody knows ?

OpenStudy (amistre64):

nobody knows what you are talking about; this is a site to help people Study and that is what all the posts here have done.

OpenStudy (amistre64):

many have even given you the formula to find a midpoint, yet you keep saying you dont need that, that you need a formula ....

OpenStudy (aravindg):

@DeviBunny try to understand what others have just told you, then you will understand that you are sitting on top of the answer!!

OpenStudy (anonymous):

Well , since nobody understands what's being asked , I'll close this and ask someone else . ; @amistre64 Yes I got the formula for a midpoint , but as I've said several times , I am not asking for the MIDPOINT formula . Please Re-read the question before you accuse . Thanks all anyways , but again , no thanks .

OpenStudy (aravindg):

@DeviBunny are you good at basic algebra?

OpenStudy (amistre64):

its not an accusation, you said: " how do I do this with a mathematical equation ? " Do you somehow define an equation and a formula differently?

OpenStudy (amistre64):

... yeah, im the one with the attitude :/ ok

OpenStudy (amistre64):

good luck